expand (1+x)^-2 using binomial theorem


1
Expert's answer
2022-03-29T17:28:32-0400

Binomial theorem

"(1+x)^{n}=1+nx+\frac{n(n-1)}{2!}x^2+..."

Here "n=-2"

"(1+x)^{-2}=1-2x+\frac{-2(-2-1)}{2!}x^2+..."

"=1-2x+\frac{6}{2!}x^2+..."

"=1-2x+\frac{6}{2}x^2+..."

"=1-2x+3x^2+...."